# Dynamical phase transitions in the non-reciprocal Ising model

### v1

@@ XII.3 Stability of the static-order phase: a droplet-capture mechanism
While the fate of the swap phase is robust to the addition of reciprocal inter-species interactions, our MC simulations suggest this is not the case for the static-order phase.
First, let us look at the right hand side of Figs. 33 and 34c-d, showing the transition to static order ($R\neq 0$ and $\mathcal{L}\approx 0$) for a finite-size system in 2D and 3D, respectively. In 2D the $\tilde{J}$-transition-point between the unstable swap regime and static order increases with system size and it is hard to conclude whether static order is stable in the thermodynamic limit. In 3D, there is some indication that there is a well-defined phase transition from disorder to static order: for linear system sizes $L=40,80,160$, there is an abrupt transition from a region in which $R$ and $\mathcal{L}$ approach zero (see $1.8\lesssim\tilde{J}\lesssim 2.2$ in Fig. 34c-d) to a regime in which $R$ is finite and $\mathcal{L}\to 0$ ($\tilde{J}\gtrsim 2.2$ in Fig. 34c-d), and the transition seems to converge at $\tilde{J}\approx 2.2$. Surprisingly, the supposed phase transition occurs even though $\tilde{K}_{+}$ is lower than $\tilde{K}_{-}$ ($\tilde{K}_{+}=0.1$ while $\tilde{K}_{-}=0.3$), which is the case where one species favors alignment while the other favors anti-alignment. This begs the question: does the droplet destabilization argument sketched in Sec. IX for the fully anti-symmetric case apply here?
@@
However, because of the asymmetry between the two species imposed by $K_{+}$, the critical droplet size $\rho_{c}$ and more importantly the velocity in which droplets expand $v(\rho)$, is different for $A$-droplets and $B$-droplets. Here, $A$-spins interact more strongly with $B$-spins than the other way around, leading to smaller $\rho_{c}$ and larger $v(\rho)$ for $A$-droplets [196]. As a result, a nested $A$-droplet inside a $B$-droplet will expand faster, eventually catching up with the boundaries of the $B$-droplets. Once this happens, our simulations show that the $B$-droplet can disintegrate, and the magnetization flip it prevented. Movie 5 shows a 2D simulation in which this mechanism takes place. Snapshots from the simulation are shown in Fig. 35.
Fig. 35: Droplet capture mechanism. MC simulation snapshots of a 2D system with $\tilde{J}=3.3$, $\tilde{K}_{+}=0.1$, $\tilde{K}_{-}=0.3$, and $L=150$. Each row displays a distinct time step, with time ascending from top to bottom. Species $B$ nucleates a droplet, but before it reaches system size, a nested droplet of $A$-spins nucleates and expands at faster speed, eventually catching up with the boundaries of the $B$-droplet, making it shrink and disappear. This mechanism can stabilize the static-order phase when $\tilde{K}_{+}<\tilde{K}_{-}$.
To summarize, when inter-species interactions are not fully anti-symmetric, there is a competing mechanism that prevents droplets from growing. This suggests a re-stabilization of the static-order phase in the general asymmetric case, for $d\geq 2$ (even if the numerical evidence from the order parameter plots is not conclusive).
@@ XII.4 Absence of infinite-period phase transition
XII.5 Schematic phase diagrams in the general asymmetric case
A schematic drawing of the 2D and 3D phase diagram with non-zero $\tilde{K}_{+}$ is shown in Fig. 36. In 2D, we only find disorder and static-order phases, while in 3D, we observe all the phases predicted by mean-field: disorder, swap, and static order. We expect the disorder-to-swap phase transition to be characterized by the same critical exponents as in Sec. VIII.1 (compatible with 3D XY), at least throughout a portion of the phase transition line. Other portions may be described by first-order-like phase transitions. Similarly, we hypothesize that the phase transition from disorder to static order in both 2D and 3D has Ising critical exponents in the regime where $K_{+}>K_{-}$, but might be first-order-like in the opposite $K_{+}<K_{-}$ regime (see Fig. 34 in which the $K_{+}<K_{-}$ phase transition seems more abrupt than the $K_{+}>K_{-}$ phase transition). Finally, we cannot exclude that the three phases in 3D meet at a tricritical point or line in the three-dimensional phase diagram.
Fig. 36: Schematic phase diagram of the general asymmetric case in (a) 2D and (b) 3D, shown as a function of $\tilde{J}$ and $\tilde{K}_{-}$, with fixed (non-zero) $\tilde{K}_{+}$. The nature of the phase transition lines are left unspecified, see discussion in Sec. XII.5. The question mark represents the uncertainty about whether the swap and static order phases meet at a point or are strictly separated by disorder.
@@ XIII Conclusions
We introduced a non-reciprocal Ising model as a minimal model for non-reciprocal many-body dynamics, and studied the stability of its static and time-dependent phases as well as the transitions between them. Our main result is the stability of a time-dependent (swap) phase in 3D, which behaves as a time-crystal with infinite coherence time in the thermodynamic limit. The phase transition from disorder to swap is characterized by critical exponents compatible with the 3D XY model, which can be rationalized by the breaking of a continuous symmetry (time translation) in the swap phase. The existence of a swap phase in 3D is robust, valid for fully anti-symmetric interactions between the species as well as in the more general case of asymmetric interactions.
Via simulations and theoretical arguments, we conjecture the static-order (ferromagnetic) phase to be unstable in any finite dimension when interactions are fully anti-symmetric. The instability is caused by the expansion of droplets of opposite magnetization occurring in the metastable species. When the inter-species interaction is not fully anti-symmetric, a competing mechanism of droplet-capture by the other species can re-stabilize the static-order phase. A direct phase transition between a time-dependent phase and a static-ordered one appears to be unfeasible.
@@ Acknowledgements.
Appendix A The effect of different update rules

### v2

@@ XII.3 Stability of the static-order phase: a droplet-capture mechanism
While the fate of the swap phase is robust to the addition of reciprocal inter-species interactions, our Monte-Carlo simulations suggest this is not the case for the static-order phase.
First, let us examine the right part of each panel of Fig. 31, showing the transition to static order ($R\neq 0$ and $\mathcal{L}\approx 0$) for a finite-size system in 2D. Figure 31a-b shows that the $\tilde{J}$-transition-point between the unstable swap regime and static order increases with system size and it is hard to conclude whether static order is stable in the thermodynamic limit. In Fig. 31c-d, where $\tilde{K}_{+}$ is varied, $R$ seems to converge to finite non-zero values (while $\mathcal{L}$ approaches zero) for $\tilde{K}_{+}\gtrsim 0.24$, which might indicate a stable phase transition from disorder to static order. Surprisingly, the supposed phase transition occurs even though $\tilde{K}_{+}$ is lower than $\tilde{K}_{-}$ ($\tilde{K}_{-}=0.3$), which is the case where one species favors alignment while the other favors anti-alignment. In 3D, Fig. 33 gives an even stronger indication that there is a well-defined phase transition from disorder to static order: for linear system sizes $L=40,80,160$, there is an abrupt transition from a region in which $R$ and $\mathcal{L}$ approach zero (see $1.8\lesssim\tilde{J}\lesssim 2.2$ in Fig. 33) to a regime in which $R$ is finite and $\mathcal{L}\to 0$ ($\tilde{J}\gtrsim 2.2$ in Fig. 33), and the transition seems to converge at $\tilde{J}\approx 2.2$. Again, the transition occurs when $\tilde{K}_{+}<\tilde{K}_{-}$ ($\tilde{K}_{+}=0.1$ while $\tilde{K}_{-}=0.3$). These indications raise the question: does the droplet destabilization argument sketched in Sec. IX for the fully anti-symmetric case apply here?
@@
However, because of the asymmetry between the two species imposed by $K_{+}$, the critical droplet size $\rho_{c}$ and more importantly the velocity in which droplets expand $v(\rho)$, is different for $A$-droplets and $B$-droplets. Here, $A$-spins interact more strongly with $B$-spins than the other way around, leading to smaller $\rho_{c}$ and larger $v(\rho)$ for $A$-droplets [233]. As a result, a nested $A$-droplet inside a $B$-droplet will expand more quickly and may eventually reach the boundaries of the $B$-droplet. Once this happens, our simulations show that the $B$-droplet can disintegrate, and the magnetization flip it prevented. Movie 5 shows a 2D simulation in which this mechanism takes place. Snapshots from the simulation are shown in Fig. 34. Note that the simulation parameters were chosen to enhance the visibility of the mechanism, allowing the $B$-droplet to persist for an extended period, but the mechanism is expected to be more efficient in suppressing droplet growth as the asymmetry between the species increases (larger $K_{+}$).
Fig. 34: Droplet capture mechanism. MC simulation snapshots of a 2D system with $\tilde{J}=3.3$, $\tilde{K}_{+}=0.1$, $\tilde{K}_{-}=0.3$, and $L=150$. See Movie 5 for the full evolution. Each row displays a distinct time step, with time ascending from top to bottom. Species $B$ nucleates a droplet, but before it reaches system size, a nested droplet of $A$-spins nucleates and expands at faster speed, eventually catching up with the boundaries of the $B$-droplet, making it shrink and disappear. This mechanism can stabilize the static-order phase when $\tilde{K}_{+}<\tilde{K}_{-}$.
To summarize, when inter-species interactions have opposite signs (the $\tilde{K}_{+}<\tilde{K}_{-}$ case) but are not fully anti-symmetric, there is a competing mechanism that prevents droplets from growing. This suggests a re-stabilization of the static-order phase in the general asymmetric case, for $d\geq 2$.
@@ XII.4 Absence of infinite-period phase transition
Fig. 35: Schematic phase diagram of the general asymmetric case in (a) 2D and (b) 3D, shown as a function of $\tilde{J}$ and $\tilde{K}_{-}$, with fixed (non-zero) $\tilde{K}_{+}$. The nature of the phase transition lines are left unspecified, see discussion in Sec. XII.5. The question mark represents the uncertainty regarding whether the swap and static-order phases meet at a point or are strictly separated by disorder.
XII.5 Schematic phase diagram in the general asymmetric case
A schematic drawing of the 2D and 3D phase diagram with a fixed non-zero $\tilde{K}_{+}$ is shown in Fig. 35. In 2D, our numerical results suggest that there are stable disorder and static-order phases (but no swap), while in 3D, our numerical results suggest that all the phases predicted by mean-field: disorder, swap, and static order, are stable. We expect the disorder-to-swap phase transition to be characterized by the same critical exponents as in Sec. VIII.1 (compatible with 3D XY), at least throughout a portion of the phase transition line. Other portions may be described by first-order-like phase transitions. Similarly, we hypothesize that the phase transition from disorder to static order in both 2D and 3D has Ising critical exponents in the regime where $K_{+}>K_{-}$, but might be first-order-like in the opposite $K_{+}<K_{-}$ regime (see Fig. 32a in which the $K_{+}<K_{-}$ phase transition seems more abrupt than the $K_{+}>K_{-}$ phase transition). Finally, we cannot exclude that the three phases in 3D meet at a tricritical point (or line in the three-dimensional phase diagram).
@@ XIII Conclusions
We introduced a nonreciprocal Ising model as a minimal model for nonreciprocal many-body dynamics, and studied the stability of its static and time-dependent phases along with the transitions between them. Our main result is the stability of a time-dependent (swap) phase in 3D, which behaves as a time-crystal with infinite coherence time in the thermodynamic limit. The phase transition from disorder to swap is characterized by critical exponents compatible with the 3D XY model, which can be rationalized by the breaking of a continuous symmetry (time translation) in the swap phase. The existence of a swap phase in 3D is robust and holds for fully anti-symmetric interactions between the species as well as the more general asymmetric interactions case.
Based on simulations and theoretical arguments, we predict that the static-order (ferromagnetic) phase is unstable in any finite dimension for fully anti-symmetric interactions. The instability is caused by the expansion of droplets of opposite magnetization occurring in the metastable species. When the inter-species interaction is not fully anti-symmetric, a competing mechanism of droplet-capture by the other species can re-stabilize the static-order phase. A direct phase transition between a time-dependent phase and a static-ordered phase, however, was not observed.
Code availability — The code used for performing the Monte-Carlo simulations, computing the critical exponents, and plotting the mean-field phase portrait is available under the 2-clause BSD license at https://doi.org/10.5281/zenodo.14816551
@@ Acknowledgements.
Appendix A Measurements of time-reversal symmetry breaking
As shown in Sec. IV.2, nonreciprocal coupling breaks detailed balance, and therefore time-reversal symmetry. This is a statement about the dynamics of our system. The dynamics can be related to thermodynamic observables – such as heat, work, and entropy – under the local detailed balance condition [156, 236, 157, 158, 237, 238]. As we don’t explicitly model the environment of our system, the local detailed balance condition is a hypothesis. Specifically, local detailed balance relates the entropy produced by a state transition (here, a spin flip) to the transition rates $w(F_{i}^{\alpha}\vec{\sigma}|\vec{\sigma})$ via
$\Delta S_{\text{single flip}}=\ln\frac{w\left(F_{i}^{\alpha}\vec{\sigma}|\vec{\sigma}\right)}{w\left(\vec{\sigma}|F_{i}^{\alpha}\vec{\sigma}\right)}.$ (75)
Here entropy is given in units of $k_{B}$. By summing over a series of transitions in a specified trajectory, the above expression gives the total entropy produced by that trajectory. For example, Eq. (17) gives the entropy produced by a specific cycle of spin states.
To characterize the steady states of the system, it is more convenient to measure the average entropy production rate. This is given by (see §6.2.2 of [156] for derivation)
$\dot{S}=\sum_{i,\alpha}\left\langle w\left(F_{i}^{\alpha}\vec{\sigma}|\vec{\sigma}\right)\ln\frac{w\left(F_{i}^{\alpha}\vec{\sigma}|\vec{\sigma}\right)}{w\left(\vec{\sigma}|F_{i}^{\alpha}\vec{\sigma}\right)}\right\rangle_{t,\Omega},$ (76)
where $\langle\ldots\rangle_{t,\Omega}$ is an average over time and realizations. Using the Glauber transition rate, Eq. (8), we obtain that the entropy production rate per spin is
$\dot{s}=-\frac{1}{L^{d}}\sum_{i,\alpha}\left\langle\left[1-\tanh\left(\frac{\Delta E^{\alpha}_{i}}{2k_{B}T}\right)\right]\frac{\Delta E_{i}^{\alpha}}{2k_{B}T}\right\rangle_{t,\Omega}.$ (77)
where $\tau$ was taken as 1 in simulation time.
The entropy production rate in Eq. (77) is a microscopic measurement of time-reversal symmetry breaking. By contrast, the phase space angular momentum $\mathcal{L}=\langle M_{B}\partial_{t}M_{A}-M_{A}\partial_{t}M_{B}\rangle_{t,\Omega}$ introduced in Eq. (VI) of the main text is a macroscopic measurement of time-reversal symmetry breaking. Coarse-graining from micro- to macroscopic variables is known to give a lower bound to the true entropy production rate [156], below we will show that $\mathcal{L}$ is proportional to the coarse-grained (mesoscopic) entropy production rate, denoted by $\dot{\mathcal{S}}$, only for cases where the fluctuations in magnetization are negligible compared with the average magnetization and if one considers the dynamics of the mean-field equation with added noise in the $m_{\alpha}\ll 1$ limit.
To see this, we work with the mean-field equation of the fully anti-symmetric case expanded around $m_{\alpha}\ll 1$, Eq. (30), as a stochastic PDE with white noise,
$-\frac{1}{3}\left(\tilde{J}m_{\alpha}+\tilde{K}\varepsilon_{\alpha\beta}m_{\beta}\right)^{3}+\eta_{\alpha}.$
where the white noise field satisfies $\langle\eta_{\alpha}(\vec{r},t)\rangle=0$ and $\langle\eta_{\alpha}(\vec{r},t)\eta_{\beta}(\vec{r}\,^{\prime},t^{\prime})\rangle=\zeta\delta_{\alpha\beta}\delta(\vec{r}-\vec{r}\,^{\prime})\delta(t-t^{\prime})$. This can be rewritten as
$\partial_{t}m_{\alpha}=-\frac{\delta\mathcal{F}}{\delta m_{\alpha}}+A_{\alpha}+\eta_{\alpha},$ (79)
where $-\delta\mathcal{F}/\delta m_{a}$ are conservative forces derived by the variation of a free energy
$\mathcal{F}=\int{\rm d}\vec{r}\bigg{[}$ $\frac{(1-\tilde{J})}{2}\left(m_{A}^{2}+m_{B}^{2}\right)+\frac{\tilde{J}^{3}}{12}\left(m_{A}^{4}+m_{B}^{4}\right)$ (80)
$+\frac{1}{2}\tilde{J}\tilde{K}^{2}m_{A}^{2}m_{B}^{2}+\frac{D}{2}\left(\left|\nabla m_{A}\right|^{2}+\left|\nabla m_{B}\right|^{2}\right)\bigg{]},$
and $A_{a}$ are the non-conservative forces on the right-hand side of Eq. (78),
$A_{\alpha}$ $=\sum_{\beta}\varepsilon_{\alpha\beta}a_{\alpha\beta}$ (81)
$a_{\alpha\beta}$ $=\tilde{K}m_{\beta}-\tilde{J}^{2}\tilde{K}m_{\alpha}^{2}m_{\beta}-\frac{1}{3}\tilde{K}^{3}m_{\beta}^{3}\,\,.$ (82)
Finally, we introduce the path probability functional of observing a trajectory $\mathbf{m}(\vec{r},t)$ as
$\mathcal{P}[\mathbf{m}(\vec{r},t)]=\frac{\exp\left(-\mathcal{A}[\mathbf{m}(\vec{r},t)]\right)}{\mathcal{Z}},$ (83)
where $\mathbf{m}\equiv(m_{A},m_{B})$, $\mathcal{Z}$ is a normalization constant, and $\mathcal{A}[\mathbf{m}(\vec{r},t)]$ is the action, which can be found using path-integral techniques [229, 239, 240, 241, 242] to be
$\mathcal{A}[\mathbf{m}(\vec{r},t)]=\dfrac{1}{\zeta}\int{\rm d}\vec{r}{\rm d}t\sum_{\alpha}\left(\partial_{t}m_{\alpha}+\dfrac{\delta\mathcal{F}}{\delta m_{\alpha}}-A_{\alpha}\right)^{2}.$ (84)
The average entropy production rate for a trajectory is written in terms of $\mathcal{P}$ and of the probability $\widetilde{\mathcal{P}}$ of observing the trajectory under time-reversal as [243]
$\dot{\mathcal{S}}=\lim_{T\to\infty}\dfrac{1}{T}\left\langle\ln\dfrac{\mathcal{P}}{\widetilde{\mathcal{P}}}\right\rangle_{\Omega}=\lim_{T\to\infty}\dfrac{\langle\widetilde{\mathcal{A}}-\mathcal{A}\rangle_{\Omega}}{T}$ (85)
where $T$ here is the trajectory duration and the average is over noise realizations [244]. As we do not have magnetic forces, finding $\widetilde{\mathcal{A}}$ amounts to taking $t\to-t$, which only affects the time derivative terms in Eq. (84), giving
$\widetilde{\mathcal{A}}[\mathbf{m}(\vec{r},t)]=\dfrac{1}{\zeta}\int{\rm d}\vec{r}{\rm d}t\sum_{\alpha}\left(-\partial_{t}m_{\alpha}+\dfrac{\delta\mathcal{F}}{\delta m_{\alpha}}-A_{\alpha}\right)^{2}.$ (86)
Plugging Eqs. (84) and (86) into Eq. (85), we find
$\dot{\mathcal{S}}$ $=\lim_{T\to\infty}\frac{4}{\zeta T}\left\langle\int{\rm d}\vec{r}{\rm d}t\sum_{\alpha}\left(A_{\alpha}-\dfrac{\delta\mathcal{F}}{\delta m_{\alpha}}\right)\partial_{t}m_{\alpha}\right\rangle_{\Omega}$ (87)
$=\lim_{T\to\infty}\frac{4}{\zeta T}\left\langle\int{\rm d}\vec{r}{\rm d}t\sum_{\alpha}\left(A_{\alpha}\partial_{t}m_{\alpha}\right)-\Delta\mathcal{F}\right\rangle_{\Omega}$
$=\lim_{T\to\infty}\frac{4}{\zeta T}\left\langle\int{\rm d}\vec{r}{\rm d}t\sum_{\alpha}A_{\alpha}\partial_{t}m_{\alpha}\right\rangle_{\Omega}$
$=\lim_{t\to\infty}\frac{4}{\zeta T}\left\langle\int{\rm d}\vec{r}{\rm d}t\left(a_{AB}\partial_{t}m_{A}-a_{BA}\partial_{t}m_{B}\right)\right\rangle_{\Omega}$
$=\frac{4}{\zeta}\int{\rm d}\vec{r}\left\langle a_{AB}\partial_{t}m_{A}-a_{BA}\partial_{t}m_{B}\right\rangle_{t,\Omega}.$
To go from the first to second line, we identified the total time derivative of the free energy, ${\rm d}\mathcal{F}/{\rm d}t=\int{\rm d}\vec{r}\sum_{\alpha}\partial_{t}m_{\alpha}\ (\delta\mathcal{F}/\delta m_{\alpha})$. To go from the second to the third line, we used the fact that the $\Delta\mathcal{F}$ is a constant, which will contribute zero to $\dot{\mathcal{S}}$ when the limit $T\to\infty$ is taken. In the last equality we replaced $\frac{1}{T}\int{\rm d}t\,...$ with a time average $\langle...\rangle_{t}$. For small $m_{\alpha}$, non-linear terms can be dropped and $a_{\alpha\beta}\simeq\tilde{K}m_{\beta}$, leading to
$\dot{\mathcal{S}}=\frac{4\tilde{K}}{\zeta}\int{\rm d}\vec{r}\langle m_{B}\partial_{t}m_{A}-m_{A}\partial_{t}m_{B}\rangle_{t,\Omega}.$ (88)
Finally, we assume small fluctuations around the mean magnetization, i.e. $m_{\alpha}(\vec{r})=M_{\alpha}+\delta m_{\alpha}(\vec{r})$, $\delta m_{\alpha}(\vec{r})\ll M_{\alpha}$, and obtain to leading order,
$\!\!\!\dot{\mathcal{S}}=\dfrac{4\tilde{K}L^{d}}{\zeta}\left\langle M_{B}\partial_{t}M_{A}-M_{A}\partial_{t}M_{B}\right\rangle_{t,\Omega}=\dfrac{4\tilde{K}L^{d}}{\zeta}\mathcal{L}.$ (89)
In conclusion, $\mathcal{L}\propto\dot{\mathcal{S}}$ for a mesoscale system obeying Eq. (78) if (i) $|m_{\alpha}|\ll 1$ and (ii) $\delta m_{\alpha}(\vec{r})\ll M_{\alpha}$. Under these two approximations, and in the vicinity of the Hopf bifurcation, Eq. (78) can be reduced into the noisy Stuart-Landau equation [85]
$\frac{{\rm d}\mathcal{M}}{{\rm d}t}=(a-b|\mathcal{M}|^{2})\mathcal{M}+H$ (90)
where $\mathcal{M}=M_{A}+iM_{B}$, $H=\eta_{A}+i\eta_{B}$, and $a$ and $b$ are complex coefficients. Writing the Stuart-Landau equation in polar coordinates and eliminating the radial equation yields a drift-diffusion equation that saturates the bound on the entropy production per oscillation $\Delta S\geq 4\pi^{2}\mathcal{N}$ of Ref. [131] mentioned in Sec. VIII.2 (here, $\mathcal{N}\equiv\tau_{c}/T_{\rm osc}$ where $\tau_{\text{c}}$ is the coherence time and $T_{\text{osc}}$ is the period of oscillation). Indeed, it is shown in Ref. [245, SM § III] that in this case, the rate of entropy production is $\dot{S}=f^{2}\tau_{\text{c}}$ in which $f=2\pi/T_{\text{osc}}$, and $\Delta S=T_{\text{osc}}\dot{S}$ saturates the bound. We refer to Ref. [245] for more details.
While $\mathcal{L}$ can be measured in any state, in general one will have $\mathcal{L}\neq\dot{\mathcal{S}}$, and no conclusion about entropy production in terms of phase space angular momentum can be made.
We calculate the entropy production rate per spin in the 3D NR Ising model using MC simulations. Figure 36 shows $\dot{s}$ for $L=40$ and $L=80$, as a function of the coupling $\tilde{J}$. The curves overlap, which shows that $\dot{s}$ is independent of system size for large systems. $\dot{s}$ decreases monotonically with $\tilde{J}$, as stronger nearest-neighbor coupling slows down the cycle rate of individual spin-pairs. By contrast, $\mathcal{L}$ is non-zero only in the swap phase. For $\tilde{J}=1.6$ and $\tilde{K}=0.1$, we compare $\dot{s}$ to the lower-bound given by the temporal autocorrelation function in Eq. (67), where $T_{\rm osc}$ is found by measuring the average oscillation time and $\tau_{c}$ is given by the fit in Fig. 17. Indeed, $\dot{s}$ exceeds the lower bound by a factor of about 10.
Fig. 36: Measurements of time-reversal symmetry breaking in the 3D NR Ising model. Entropy production rate per spin, $\dot{s}$ (blue and orange squares, calculated from Eq. (77)), and phase space angular momentum, $\mathcal{L}$ (green dots), as a function of $\tilde{J}$ for fixed $\tilde{K}=0.1$. Theoretical lower bound for $\dot{s}$, $\dot{s}^{*}=\frac{4\pi^{2}}{T_{\rm osc}^{2}}\,\frac{\tau_{c}(L)}{L^{d}}$ (Eq. (67)), at $\tilde{J}=1.6$, is shown as a red star, where $T_{\rm osc}$ was measured to be $T_{\rm osc}=120$ and $\tau_{c}/L^{3}$ was taken as $0.32$ in accordance with the fit in Fig. 17. $\dot{s}^{*}=9\times 10^{-4}$ while $\dot{s}=0.01$ at $\tilde{J}=1.6$.
Appendix B 1D NR Ising model
Fig. 37: 1D NR Ising model. (a) $R$ and (b) $\mathcal{L}$ as a function of $\tilde{J}$ for $\tilde{K}=0.3$ and different linear system size $L$. (c) Kymographs of $\theta$ for $\tilde{J}=3$ (left) and $\tilde{J}=4$ (right), comparing the Ising model ($\tilde{K}=0$, top) with the NR Ising model ($\tilde{K}=0.2$, middle and $\tilde{K}=0.5$, bottom). System size is $L=500$. Obtained from MC simulations.
The 1D NR Ising model does not have a phase transition to an ordered state (either static or time dependent), as is evident from the decrease of $R$ and $\mathcal{L}$ as system size increases. Similarly to the 1D Ising model, as $L$ increases, domains of opposite magnetization emerge and the system becomes disordered. However, unlike the Ising model, the boundaries between the domains do not diffuse but move with non-zero velocity (see Fig. 37c where the kymograph slope indicates the velocity). Domains expand in a way that can be viewed as the 1D version of droplet growth discussed in Sec. IX.
Appendix C The effect of different update rules